OpenStudy (anonymous):

Fin the angle between u=<2sqrt(2), -4> and v= <-sqrt(2), 1>

OpenStudy (anonymous):

\[cos(\theta)=\frac{u\cdot v}{|u||v|}\]

OpenStudy (anonymous):

so we need these numbers. \[u\cdot v=2\sqrt{2}\times -\sqrt{2}+(-4)\times 1=-8\]

OpenStudy (anonymous):

if my arithmetic is correct, please check it

OpenStudy (anonymous):

\[|u|=\sqrt{(2\sqrt{2})^2+4^2}=\sqrt{24}\]

OpenStudy (anonymous):

\[|v|=\sqrt{(\sqrt{2})^2+1^2}=\sqrt{5}\] and so \[|u||v|=\sqrt{120}=2\sqrt{30}\]\]

OpenStudy (anonymous):

\[\cos(\theta)=\frac{-4}{\sqrt{30}}\]

OpenStudy (anonymous):

i get \[\theta = 136.9\] in degrees, but you should check my arithmetic i make misakes

OpenStudy (anonymous):

alright, thanks man

OpenStudy (anonymous):

yw, but please check, especially if you have to hand this in